# Mitigating the Stability-Plasticity Dilemma in Adaptive Train Scheduling with Curriculum-Driven Continual DQN Expansion

### v1

@@ Introduction
Continual learning agents, capable of acquiring new knowledge and skills over time, are essential for environments that constantly change (Abel et al. 2024; Khetarpal et al. 2022). One central challenge for continual learning systems is balancing the retention of previously learned knowledge with the acquisition of new knowledge—a concept known as the stability-plasticity dilemma (Schwarz et al. 2018). This dilemma is relevant in many real life application domains such as the train scheduling problem (TSP), where environmental conditions (i.e routes, schedules) and agent behaviors (i.e train characteristics) are constantly changing, and the search space is vast. TSP is inherently complex. Even in simplified scenarios, agents often require considerable time to converge on optimal solutions (Toth and Vigo 2002).
@@
One approach to address the complexity of TSP is curriculum learning, which introduces a sequence of environments that gradually increase in difficulty, allowing the agent to build upon previously acquired skills (Narvekar et al. 2020). This approach mirrors how humans learn, starting with simpler concepts and progressively tackling more complex ones. However, curriculum learning introduces its own set of challenges due to the non-stationary training environment.
@@
Various continual reinforcement learning (CRL) approaches have been proposed to mitigate catastrophic forgetting, improve plasticity, facilitate transfer learning across new tasks, and scale effectively to handle multiple tasks. Current methods can be categorized into fixed-size approaches, which do not change the size of the network for new tasks (i.e., regularization-based approaches or optimization-based methods), and network expansion approaches. While network expansion approaches are more likely to achieve better performance, they may suffer from scaling issues as the number of tasks increases (Wang et al. 2023).
@@
To address this, recent studies in the mode connectivity literature have investigated whether multitask and continual solutions are connected by a manifold of low error. Results show that different optima obtained by gradient-based optimization methods are connected by simple paths of non-increasing loss (Mirzadeh et al. 2020). Inspired by this, various works have demonstrated how to efficiently build simplicial complexes for fast ensembling and adaptively grow the subspace as the system encounters new tasks. This can be leveraged to develop neural systems that scale better while maintaining good performance across multiple tasks (Benton et al. 2021; Gaya et al. 2022).
@@
While previous works focused on continual learning with distinct tasks, a curriculum does not involve distinct tasks with clear boundaries but rather a series of adjacent skills that build on top of each other, such as pathfinding, dealing with unexpected malfunctions, or avoiding deadlocks. This introduces a unique type of non-stationarity, where agents are learning interrelated skills. The challenge lies in effectively sequencing these skills to maximize learning efficiency and performance while also addressing the stability-plasticity dilemma for the learning agents.
@@
In this work, we propose a well-structured curriculum for the train scheduling problem, decomposing it into various skills like pathfinding, malfunction handling, and deadlock avoidance. Furthermore, we investigate different approaches to leverage this curriculum and introduce the Continual Deep Q-Network Expansion (CDE) algorithm. The core idea of our CDE approach is to dynamically expand, adjust, and prune Q-function subspaces based on current task changes in the environment. The decision-making process within CDE involves determining whether a pre-existing Q-function subspace is effective for a new task and thus adapting it, or if a new subspace needs to be created for faster learning. The pre-existing Q-function subspaces are adapted to new tasks with elastic weight consolidation to avoid catastrophic forgetting. while the newly created subspaces are trained with adaptive rational activation functions for higher plasticity. In the pruning step, we retain only the best Q-networks to maintain sublinear scaling. Experimental results demonstrate significant improvements compared to other approaches, highlighting the potential of CDE to balance stability and plasticity, leading to more efficient and adaptable learning [^1] A link to the code will be added soon..
@@
Continual Learning Traditional RL methods, which assume static environments, often lead to models prone to overfitting and lacking adaptability to new, varying contexts. Continual RL (CRL) aims to develop adaptive models that learn from a continuous data stream and evolve across multiple tasks by leveraging accumulated knowledge, thereby overcoming the limitations of static-task assumptions (Abel et al. 2024).
@@
Addressing Catastrophic Forgetting: Efforts such as those by Rusu et al. (2016), Cheung et al. (2019), and Wortsman et al. (2020) aim to preserve model parameters from previously learned tasks but often do not scale well with an increasing number of tasks mainly due to substantial computational and memory requirements. Other strategies, like maintaining a memory buffer of prior experiences (Lopez-Paz and Ranzato 2017; Isele and Cosgun 2018; Caccia et al. 2020), also face scalability issues as memory costs increase with task complexity and number. Methods enhancing transfer to new tasks, such as fine-tuning, are scalable and adaptable but prone to catastrophic forgetting. Elastic weight consolidation (Kirkpatrick et al. 2017) addresses this by limiting weight modifications, though it may reduce network’s plasticity as the number of tasks increase. Knowledge distillation techniques, as discussed in (Rusu et al. 2015; Li and Hoiem 2017; Rolnick et al. 2019), and methods leveraging commonalities across tasks (Xu and Zhu 2018; Pasunuru and Bansal 2019), have been utilized to improve CRL performance. Meta-learning and generative models also enhance adaptability and learning efficiency (Javed and White 2019; Beaulieu et al. 2020; Shin et al. 2017). However, practical application challenges persist, as indicated by (Lee et al. 2020), who propose an iteratively expanding network model primarily for supervised learning without distinct task boundaries.
@@
Addressing Neural Plasticity: While addressing catastrophic forgetting is essential for non-stationary agents, extensive neural plasticity is required due to inherent distribution shifts during agent training, such as input and output drifts (Mnih et al. 2015). These shifts are exemplified by various environment dynamics as agents adapt to changing states and rewards. Recent studies have explored approaches to increase plasticity and allow faster adaptation to changing environments, such as concatenated ReLU (Abbas et al. 2023) and the introduction of new trainable weights (Nikishin et al. 2024). Additionally, techniques like continual backpropagation (Dohare et al. 2023) and dynamic hyperparameter and activation function adaptation (Loni et al. 2023; Delfosse et al. 2021) highlight the potential of dynamically alter the weight optimization landscape,which can aid continual adaptation and thus paving the way for more robust RL systems.
Curriculum Learning Curriculum learning is a learning paradigm where training data is organized in a progressively challenging manner, introducing concepts gradually to improve learning outcomes (Bengio et al. 2009). Curriculum learning is applied across various fields demonstrating improved sample efficiency and better convergence of machine learning algorithms (Soviany et al. 2022). Curriculum learning not only helps in achieving better performance on evaluation data but also in handling noisy datasets by starting with cleaner data and gradually introducing noisier parts (Wang, Chen, and Zhu 2021).
In reinforcement learning, curriculum learning has the potential to enhance learning efficiency and performance(Narvekar et al. 2020). Early applications of curriculum learning in RL were primarily manual, such as the work by Stanley and Miikkulainen (2005), where domain experts designed curricula for video game agents. Recently, there has been a shift towards automating curriculum generation. For instance, Narvekar and Stone (2019) proposed learning a curriculum policy using a Curriculum Markov Decision Process (CMDP), which maps the learning progress to the subsequent tasks, showing that intermediate tasks do not always need to be learned to convergence. Other methods, such as those by Foglino, Coletto Christakou, and Leonetti (2019) and Foglino et al. (2019), frame curriculum sequencing as a combinatorial optimization problem, employing metaheuristic algorithms like beam search and genetic algorithms to identify optimal task sequences.
Despite these advancements, several challenges persist in curriculum learning for RL. One significant issue is the computational expense associated with learning a curriculum policy independently for each agent and task, as highlighted by Narvekar and Stone (2019). Another challenge is catastrophic forgetting, where knowledge from previous tasks is lost when learning new tasks. Additionally, when state and action spaces differ between tasks, transferring high-level knowledge such as partial policies or options becomes crucial, as explored by Yang and Asada (1996) and Zimmer et al. (2018).
Train Scheduling Problem
Mathematical Formulation
@@
TSP is an NP-Hard optimization and logistics challenge focused on the optimal assignment of trains to routes and their scheduling across a railway network while adhering to diverse constraints (Caprara, Fischetti, and Toth 2002). It requires dynamic replanning to adapt to unforeseen disruptions effectively. This issue integrates aspects of Multi-Agent Pathfinding and the Vehicle Rescheduling Problem, combining approaches from both fields to enhance the efficacy of train scheduling strategies (Cai and Goh 1994).
@@
Given $K$ (number of trains), $N=\{1,\ldots,n\}$ (set of yards), $S_{k}$ (stop-schedule for train $k$), $v_{k}$ (speed of train $k$), $m_{k}$ (malfunction rate of train $k$), $h_{k}$ (time steps train $k$ is halted in case of a malfunction), $d_{ij}$ (distance between yards $i$ and $j$), $x_{ij}^{k}$ (binary variable indicating if train $k$ travels directly from yard $i$ to yard $j$), the objective is to minimize the total travel time for all trains, including delays due to malfunctions. The mathematical formulation of the TSP can be represented as follows:
@@
Traditionally, Operations Research (OR) techniques are employed to tackle the TSP, leveraging mathematical modeling and optimization to systematically improve train path scheduling (Lusby et al. 2011; Dong et al. 2020). These methods excel in breaking down the TSP into a graph optimization problem, aiming to identify non-intersecting routes for trains in time and space.
However, the reliance on OR comes with limitations, particularly in its flexibility and responsiveness to unexpected disruptions, such as weather conditions or infrastructure or train failures. Additionally, as railway networks expand and the number of operational trains increases, the computational demands of OR methods surge, leading to significant challenges in maintaining optimal schedules, especially under conditions requiring frequent rescheduling (Khadilkar 2018).
Flatland Simulator
The Flatland Simulator is two-dimensional simplified grid environment that allows for faster and diverse experimentation in TSP. Flatland does not only reduce the complexity of the full physical simulation, but also provides an easy-to-use interface to test novel approaches for the TSP (Mohanty et al. 2020).
Agents are tasked with efficiently navigating trains from start to finish, with the complexity of tasks escalating as the grid expands and more trains are added.
The score is computed based on the successful routing of trains to their destinations and the completion of scenarios, where all trains within an environment reach their respective endpoints, testing their strategies for effective train scheduling and rescheduling under varying conditions.
Curriculum Design for the Train Scheduling Problem
To construct an effective curriculum, it is essential to consider task generation, sequencing, and transfer learning. A well-designed curriculum improves both accuracy and speed by breaking down the train scheduling problem into distinct subproblems enabling better performance and faster convergence.
In designing our curriculum for the TSP, we focused on addressing four key challenges integral to efficient rail system management:
@@
  • Deadlocks: It is critical that an agent efficiently handles various types of deadlocks (common, cyclic, and track-caused; see Appendix for visualization and further details). An agent should efficiently plan and avoid paths that might lead to these scenarios, even at the cost of increased travel time for some trains.
These challenges often occur in combination, requiring agents to make decisions that minimize the impact of such scenarios on overall system efficiency.
Based on these challenges, we design a curriculum tailored to address specific train scheduling problems (TSP) by alternating between three specialized environment types focused on pathfinding, malfunctions and train-speed differences, and deadlocks (see Figure LABEL:fig:main).
@@
  • Deadlock avoidance focuses on the most common deadlock scenarios, gradually increasing complexity by adding more agents and reducing number of switches.
In this work, we sequence the tasks in the following way: Pathfinding, malfunctions and train-speed adjustments, and deadlock avoidance. We posit that starting with pathfinding is the most sensible choice, as all other tasks require an efficient pathfinding strategy. We leave deadlock avoidance for last, as it requires the agent to master the skills of pathfinding and handling malfunctions and various train speeds.
@@
Continual DQN Expansion Algorithm
We propose a novel approach for adaptively constructing Q-function subspaces to address the challenges of non-stationary environments. Our Continual Deep Q-Network (DQN) Expansion (CDE) approach dynamically generates and adjusts Q- function subspaces, denoted as $\Theta_{1},\ldots,\Theta_{N}$, based on environmental changes and task requirements. This method not only aims to mitigate catastrophic forgetting through selective expansion and the use of an elastic weight consolidation regularized loss but also seeks to enhance the agent’s plasticity by leveraging rational adaptive activation function for faster adaptability to new tasks.
In the subsequent sections, we provide a brief overview of elastic weight consolidation (EWC) and padé activation units (PAU). Following this, we explain our methodology, detailing the integration of our proposed subspace expansion technique with EWC and PAU.
Formalisation of Continual Reinforcement Learning Problem
A continual reinforcement learning (RL) scenario is characterized by a sequence of $N$ tasks or instances. Each task $k$ is described through a Markov Decision Process (MDP) $M_{k}=\langle S_{k},A_{k},T_{k},r_{k},\gamma\rangle$, which encompasses a set of states $S_{k}$, a set of actions $A_{k}$, a transition function $T_{k}:S_{k}\times A_{k}\rightarrow\mathcal{P}(S_{k})$, and a reward function $r_{k}:S_{k}\times A_{k}\rightarrow\mathbb{R}$. We assume in our case that all tasks share identical state and action spaces.

### v2

@@ 1 Introduction
Continual learning agents, capable of acquiring new knowledge and skills over time, are essential for environments that constantly change Abel et al. (2024); Khetarpal et al. (2022). One central challenge for continual learning systems is balancing the retention of previously learned knowledge with the acquisition of new knowledge—a concept known as the stability-plasticity dilemma Schwarz et al. (2018). This dilemma is relevant in many application domains such as the train scheduling problem (TSP), where environmental conditions (i.e. routes, schedules) and agent behaviors (i.e. train characteristics) are constantly changing, and the search space is vast. TSP is inherently complex. Even in simplified scenarios, agents often require considerable time to converge to optimal solutions Toth & Vigo (2002).
@@
One approach to address the complexity of TSP is curriculum learning, which introduces a sequence of environments that gradually increase in difficulty, allowing the agent to build upon previously acquired skills Narvekar et al. (2020). This approach mirrors how humans learn, starting with simpler concepts and progressively tackling more complex ones. However, curriculum learning introduces its own set of challenges regarding non-stationary learning.
@@
Various continual reinforcement learning (CRL) approaches have been proposed to mitigate catastrophic forgetting, improve plasticity, facilitate transfer learning across new tasks, and scale effectively to handle multiple tasks. Current methods can be categorized into fixed-size approaches, which do not change the size of the network for new tasks (i.e., regularization-based approaches or optimization-based methods), and network expansion approaches. While network expansion approaches are more likely to achieve better performance, they may suffer from scaling issues as the number of tasks increases Wang et al. (2023).
@@
To address this, recent studies in the mode connectivity literature have investigated whether multitask and continual solutions are connected by a manifold of low error. Results show that different optima obtained by gradient-based optimization methods are connected by simple paths of non-increasing loss Mirzadeh et al. (2020). Inspired by this, various works have demonstrated how to efficiently build simplicial complexes for fast ensembling and adaptively grow the subspace as the system encounters new tasks. This can be leveraged to develop neural systems that scale better while maintaining good performance across multiple tasks Benton et al. (2021); Gaya et al. (2022).
@@
While previous works focused on continual learning with distinct tasks, a curriculum does not involve distinct tasks with clear boundaries but rather a series of adjacent skills that build on top of each other, such as pathfinding, dealing with unexpected malfunctions, or avoiding deadlocks in TSP. This introduces a unique type of non-stationarity, where agents are learning interrelated skills. The challenge lies in effectively sequencing these skills to maximize learning efficiency and performance while also addressing the stability-plasticity dilemma for the learning agents.
@@
In this work, we propose a well-structured curriculum for the train scheduling problem, decomposing it into various skills like pathfinding, malfunction handling, and deadlock avoidance. Furthermore, we investigate different approaches to leverage this curriculum and introduce the Continual Deep Q-Network Expansion (CDE) algorithm. The core idea of our CDE approach is to dynamically expand, adjust, and prune Q-function subspaces based on current task changes in the environment. The decision-making process within CDE involves determining whether a pre-existing Q-function subspace is effective for a new task and thus adapting it, or if a new subspace needs to be created for faster learning. The pre-existing Q-function subspaces are adapted to new tasks with elastic weight consolidation to avoid catastrophic forgetting. while the newly created subspaces are trained with adaptive rational activation functions for higher plasticity. In the pruning step, we retain only the best Q-networks to maintain sublinear scaling. Experimental results demonstrate significant improvements compared to other approaches, highlighting the potential of CDE to balance stability and plasticity, leading to more efficient and adaptable learning [^1] A link to the code will be added in the camera ready version..
@@ 2 Related Work
Continual Learning Traditional RL methods, which assume static environments, often lead to models prone to overfitting and lacking adaptability to new, varying contexts. Continual RL (CRL) aims to develop adaptive models that learn from a continuous data stream and evolve across multiple tasks by leveraging accumulated knowledge, thereby overcoming the limitations of static-task assumptions Abel et al. (2024).
@@
Addressing Catastrophic Forgetting: Efforts such as those by Rusu et al. (2016), Cheung et al. (2019), and Wortsman et al. (2020) aim to preserve model parameters from previously learned tasks but often do not scale well with an increasing number of tasks mainly due to substantial computational and memory requirements. Other strategies, like maintaining a memory buffer of prior experiences Lopez-Paz & Ranzato (2017); Isele & Cosgun (2018); Caccia et al. (2020), also face scalability issues as memory costs increase with task complexity and number. Methods enhancing transfer to new tasks, such as fine-tuning, are scalable and adaptable but prone to catastrophic forgetting. Elastic weight consolidation Kirkpatrick et al. (2017) addresses this by limiting weight modifications, although it may reduce the network plasticity as the number of tasks increases. Knowledge distillation techniques, as discussed in Rusu et al. (2015); Li & Hoiem (2017); Rolnick et al. (2019), and methods leveraging similarities across tasks Xu & Zhu (2018); Pasunuru & Bansal (2019), have been utilized to improve CRL performance. Meta-learning and generative models also enhance adaptability and learning efficiency Javed & White (2019); Beaulieu et al. (2020); Shin et al. (2017). However, practical application challenges persist, as indicated by Lee et al. (2020), who propose an iteratively expanding network model primarily for supervised learning without distinct task boundaries.
@@
Addressing Neural Plasticity: While addressing catastrophic forgetting is essential for non-stationary agents, extensive neural plasticity is required due to inherent distribution shifts during agent training, such as input and output drifts Mnih et al. (2015). These shifts are exemplified by various environment dynamics as agents adapt to changing states and rewards. Recent studies have explored approaches to increase plasticity and allow faster adaptation to changing environments, such as concatenated ReLU Abbas et al. (2023) and the introduction of new trainable weights Nikishin et al. (2024). Additionally, techniques like continual backpropagation Dohare et al. (2023) and dynamic hyperparameter and activation function adaptation Loni et al. (2023); Delfosse et al. (2021) highlight the potential of dynamically altering the weight optimization landscape.
Curriculum Learning Curriculum learning arranges training data in gradually increasing difficulty to introduce concepts incrementally and improve learning outcomes Bengio et al. (2009). It has demonstrated enhanced sample efficiency and convergence across various machine learning applications Soviany et al. (2022); Wang et al. (2021). In reinforcement learning (RL), curriculum learning can boost efficiency and performance Narvekar et al. (2020). Early RL curricula were manually designed by domain experts, for instance in Stanley & Miikkulainen (2005), where video game agents were trained via expert-crafted task sequences.
Recent work has shifted toward automated curriculum generation. For example, Narvekar & Stone (2019) introduced a Curriculum Markov Decision Process that selects subsequent tasks based on learning progress. Other approaches cast curriculum sequencing as a combinatorial optimization problem, using metaheuristics like beam search or genetic algorithms to identify optimal task orders Foglino et al. (2019a; b). Florensa et al. (2017) proposed “reverse” training, starting from goal states and progressively moving to initial states farther from the goal, thus automatically adapting the curriculum to the agent’s performance. Nevertheless, learning a separate curriculum policy for each agent and task is computationally expensive Narvekar & Stone (2019), and catastrophic forgetting can occur when transitioning between tasks. Furthermore, when tasks differ in state and action spaces, transferring high-level knowledge (e.g., partial policies or options) becomes critical Yang & Asada (1996); Zimmer et al. (2018).
Figure 2: Illustration of the curriculum designed for TSP through distinct environments. (a) Pathfinding Environment: A grid where a single agent navigates from start to destination, emphasizing efficient route planning. (b) Malfunctions and Train Speeds Environment: A setting where agents deal with random malfunctions and varying train speeds, requiring adaptive responses and strategic planning to minimize delays. (c) Deadlocks Environment: Scenarios focusing on common deadlock situations, gradually increasing in complexity with more agents and fewer switches, to train agents in avoiding gridlock. (d) Full Task Evaluation Environment: An integrated setting combining pathfinding, malfunctions, train speeds, and deadlock scenarios to evaluate the agents. The curriculum environments (a-c) are available in multiple sizes to provide diverse training scenarios, as detailed in the appendix. This curriculum structure improves learning efficiency and performance on the full task evaluation environment (d).
3 Train Scheduling Problem
3.1 Mathematical Formulation
@@
TSP is an NP-Hard optimization problem focused on the optimal assignment of trains to routes and their scheduling across a railway network while adhering to diverse constraints Caprara et al. (2002). It requires dynamic replanning to adapt to unforeseen disruptions. This issue integrates aspects of Multi-Agent Pathfinding and the Vehicle Rescheduling Problem Cai & Goh (1994).
@@
Given $K$ (number of trains), $N=\{1,\ldots,n\}$ (set of yards), $S_{k}$ (stop-schedule for train $k$), $v_{k}$ (speed of train $k$), $m_{k}$ (malfunction rate of train $k$), $h_{k}$ (time steps train $k$ is halted in case of a malfunction), $d_{ij}$ (distance between yards $i$ and $j$), $x_{ij}^{k}$ (binary variable indicating if train $k$ travels directly from yard $i$ to yard $j$), the objective is to minimize the total travel time for all trains, including delays due to malfunctions. The mathematical formulation is as follows:
@@
Traditionally, Operations Research (OR) techniques are employed to tackle the TSP, leveraging mathematical modeling and optimization to systematically improve train path scheduling Lusby et al. (2011); Dong et al. (2020). These methods break down the TSP into a graph optimization problem to identify non-intersecting routes for trains in time and space. While OR methods provide structured solutions, they face notable limitations, particularly in computational efficiency and adaptability. Computing exact solutions incurs high overhead, and frequent rescheduling is required in this domain to accommodate disruptions such as infrastructure failures, or train delays Khadilkar (2018).
3.2 Flatland Simulator
The Flatland Simulator is a two-dimensional simplified grid environment that allows for faster and diverse experimentation in TSP. Flatland provides an easy-to-use interface to test novel approaches for the TSP Mohanty et al. (2020). Agents are tasked with efficiently navigating trains from start to finish. The score is computed based on the successful routing of trains to their destinations and the completion of scenarios, where all trains within an environment reach their respective endpoints.
4 Curriculum Design for the Train Scheduling Problem
To construct an effective curriculum, it is essential to consider task generation, sequencing, and transfer learning. A well-designed curriculum improves both accuracy and speed by breaking down the train scheduling problem into distinct subproblems enabling better performance and faster convergence. Our approach specifically targets four key challenges essential for TSP:
@@ 4 Curriculum Design for the Train Scheduling Problem
  • Deadlocks: It is critical that an agent efficiently handles various types of deadlocks (common, cyclic, and track-caused). An agent should efficiently plan and avoid paths that might lead to these scenarios, even at the cost of increased travel time for some trains.
These challenges often occur in combination, requiring agents to make decisions that minimize the impact of such scenarios on overall system efficiency. We design a curriculum tailored to address specific problems by alternating between three specialized environment types focused on pathfinding, malfunctions and train-speed differences, and deadlocks (see Figure 2).
@@
  • Deadlock avoidance focuses on the most common deadlock scenarios, where two or more trains become mutually blocked at junctions or intersections and preventing any movement.
We sequence the tasks in the following way: Pathfinding, malfunctions and train-speed adjustments, and deadlock avoidance. We posit that starting with pathfinding is the most sensible choice, as all other tasks require an efficient pathfinding strategy. We leave deadlock avoidance for last, as it requires the agent to master the skills of pathfinding and handling malfunctions and various train speeds.
@@
5 Continual DQN Expansion Algorithm
We propose a novel approach for adaptively constructing Q-function subspaces to address the challenges of non-stationary environments. Our Continual Deep Q-Network (DQN) Expansion (CDE) approach dynamically generates and adjusts Q- function subspaces, denoted as $\Theta_{1},\ldots,\Theta_{N}$, based on environmental changes and task requirements. By selectively expanding these subspaces and employing elastic weight consolidation (EWC), our method mitigates catastrophic forgetting, while a rational adaptive activation function enhances plasticity for rapid task adaptation. In the following sections, we review EWC and Padé activation units (PAU) and describe how our methodology integrates subspace expansion with PAU and EWC.
5.1 Formalisation of the Continual Reinforcement Learning Problem
A continual reinforcement learning scenario is characterized by a sequence of $N$ tasks or instances. Each task $k$ is described through a Markov Decision Process (MDP) $M_{k}=\langle S_{k},A_{k},T_{k},r_{k},\gamma\rangle$, which encompasses a set of states $S_{k}$, a set of actions $A_{k}$, a transition function $T_{k}:S_{k}\times A_{k}\rightarrow\mathcal{P}(S_{k})$, and a reward function $r_{k}:S_{k}\times A_{k}\rightarrow\mathbb{R}$. We assume in our case that all tasks share identical state and action spaces.